OpenStudy (anonymous):

Remove Parentheses and Simplify 6q+6q-3(6q-9q+5r)

OpenStudy (phi):

multiply the -3 times each term inside the parentheses to get rid of the parens.

jhonyy9 (jhonyy9):

12q -18q +27q -15r =21q -15r =3(7q-5r)